OpenStudy (anonymous):

Use guess-and-check to find two numbers that have the sum of 10 and a difference of 6.

OpenStudy (anonymous):

10=x-6 so then 10=6+4

OpenStudy (anonymous):

x+y=10 x-y=6 SOLVE FOR x and y

OpenStudy (anonymous):

It's guess and check or most commonly called "trial and error", you're not meant to solve it as a system of equations.

OpenStudy (anonymous):

But then again, that's for future reference when doing questions without trial and error.